# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ Abstract
We introduce an annotated extension of treewidth that measures the contribution of a vertex set $X$ to the treewidth of a graph $G.$ This notion provides a graph distance measure to some graph property $\mathcal{P}$: A vertex set $X$ is a $k$-treewidth modulator of $G$ to $\mathcal{P}$ if the treewidth of $X$ in $G$ is at most $k$ and its removal gives a graph in $\mathcal{P}.$ This notion allows for a version of the Graph Minors Structure Theorem (GMST) that has no need for apices and vortices: $K_{k}$-minor free graphs are those that admit tree-decompositions whose torsos have $c_{k}$-treewidth modulators to some surface of Euler-genus $c_{k}.$ This reveals that minor-exclusion is essentially tree-decomposability to a “modulator-target scheme” where the modulator is measured by its treewidth and the target is surface embeddability.
We then fix the target condition by demanding that $\Sigma$ is some particular surface and define a “surface extension” of treewidth, where $\Sigma\mbox{-}{\sf tw}(G)$ is the minimum $k$ for which $G$ admits a tree-decomposition whose torsos have a $k$-treewidth modulator to being embeddable in $\Sigma.$ We identify a finite collection $\mathfrak{D}_{\Sigma}$ of parametric graphs and prove that the minor-exclusion of the graphs in $\mathfrak{D}_{\Sigma}$ precisely determines the asymptotic behavior of ${\Sigma}\mbox{-}{\sf tw},$ for every surface $\Sigma.$ It follows that the collection $\mathfrak{D}_{\Sigma}$ bijectively corresponds to the “surface obstructions” for $\Sigma,$ i.e., surfaces that are minimally non-contained in $\Sigma.$
Keywords: Treewidth, Graph minor, Surface embedding, Graph modulator, Graph Genus, Parametric graph, Universal obstruction, Graph parameter, Graph structure theorem.
@@ 1 Introduction
An underlying fact in many results in structural graph theory is that graphs that do not “contain” some graph can be seen as the result of gluing together simpler graphs of some particular property in a tree-like fashion. When the containment relation is the minor relation [^1] We say that a graph $H$ is a minor of a graph $G$ if $H$ can be obtained from some subgraph of $G$ by contracting edges. this property is typically enjoying certain topological characteristics. Moreover, in some cases, such structural theorems may serve as the combinatorial base for algorithmic applications.
Two celebrated, and highly related, theorems of this type are the Grid Theorem and the Graph Minors Structure Theorem (GMST), proven by Robertson and Seymour in [39] and  [42] respectively.
The Grid Theorem asserts that every graph which cannot be obtained by gluing together small graphs in a tree-like fashion must contain all planar graphs up to a certain size as minors. Vice-versa, if some graph $G$ does not contain a fixed planar graph $H$ as a minor, $G$ can indeed be obtained by gluing together small graphs in a tree-like fashion where the notion of “small” now depends on the excluded graph $H.$ Taking this insight as the base case, the GMST gives the most general version of similar statement. Here, the excluded minor is no longer planar, but a clique of fixed size. Any graph excluding such a large clique minor can be obtained by gluing together graphs that “almost” embed in some surface of bounded Euler-genus in, again, a tree-like fashion. This is a property that is not enjoyed by large complete graphs.
The name “Grid Theorem” stems from the way Robertson and Seymour prove the above result: They identify a single parametric graph, namely the $(n\times n)$-grid, and show that
@@
    every planar graph is a minor of the $(n\times n)$-grid for some $n\in\mathbb{N},$
@@
    the treewidth of the $(n\times n)$-grid depends on $n,$ and
  3. 3.
    there exists a function $f$ such that for every $n\in\mathbb{N},$ every graph with treewidth more than $f(n)$ contains the $(n\times n)$-grid as a minor.
So, the grid fulfils two roles: It acts as a universal obstruction for the parameter treewidth and it is an omnivore for the class of planar graphs; that is, the class of all minors of some grid is exactly the class of planar graphs. This means that embeddability on the sphere is tightly captured by the decomposability into small pieces, defined by the parameter of treewidth.
This paper is motivated by the following question:
@@ 1.1 Our contribution
Our strategy to resolve the motivating question above follows the three ingredients given for treewidth. We identify families of grid-like graphs, we call them Dyck-grids (or surface-grids if we are talking about their less streamlined version). Each Dyck-grid is a parametric graph which is an omnivore for all graphs that embed in the respective surface. We then proceed by proving that the minor exclusion of the Dyck-grids corresponding to some set of surfaces $\mathbb{S}$ leads to a “surface” generalization of treewidth, namely ${\mathbb{S}}\mbox{-}{\sf tw},$ defined via tree-decompositions where all torsos “almost embed” into some surface in $\mathbb{S}.$ Finally, we prove that, with respect to our notion of “almost embedability”, the Dyck-grids themselves cannot be decomposed in this way. With this, we obtain an asymptotic characterization of the resulting width parameters ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}$ in form of the exclusion of parametric graphs.
Measuring modification by treewidth.
To capture what we mean by “almost embeddable” we use again treewidth. In particular, we introduce an annotated extension of treewidth as follows.
Given a graph $H$ and a set $X\subseteq V(G),$ we say that $H$ is an $X$-rooted minor of $G$ (or, simply, an $X$-minor of $G$) if there is a collection $\mathcal{S}=\{S_{v}\mid v\in V(H)\}$ of pairwise vertex-disjoint connected [^2] A set $X\subseteq V(G)$ is connected in $G$ if the induced subgraph $G[X]$ is a connected graph. subsets of $V(G),$ each containing at least one vertex of $X$ and such that, for every edge $xy\in E(H),$ the set $S_{x}\cup S_{y}$ is connected in $G.$ $H$ is a minor of $G,$ denoted by $H\leq G,$ if $H$ is a $V(G)$-minor of $G.$
Given an annotated graph $(G,X)$ where $G$ is a graph and $X\subseteq V(G),$ we define ${\sf tw}(G,X)$ as the maximum treewidth of an $X$-minor of $G.$
Notice that the annotated extension of treewidth above can be defined for every parameter (see (12) in Section 7). In the case of treewidth, this gives a measure of the contribution of $X$ to the tree decomposability of $G.$ We say that an $X\subseteq V(G)$ is a $k$-treewidth modulator of $G$ to some graph class $\mathcal{G}$ if ${\sf tw}(G,X)\leq k$ and $G-X\in\mathcal{G}.$ That way, treewidth may serve as the measure of the amount of modification caused by the removal of a modulator to some target graph class. In the case of our results, the target graph class will be the class $\mathcal{E}_{g}$ containing all graphs of Euler-genus at most $g.$
As a first step, we display how the GMST can be restated using the modulator-target duality above. For this, we consider the Hadwiger number of a graph, that is the graph parameter ${\sf hw},$ where ${\sf hw}(G)$ is the largest size of a clique minor of $G.$ We define the parameter ${\sf hw}^{\prime}$ as follows
The clique-sum closure, defined formally in Section 2, formalizes the “tree-decomposability” notion that underlines structural theorems such as the Grid Theorem and the GMST.
In this paper we provide a variant of the GMST stating that ${\sf hw}$ and ${\sf hw}^{\prime}$ are asymptotically equivalent (Theorem 8.2). That is, there exists a function $f\colon\mathbb{N}\to\mathbb{N}$ such that for all graphs $G$ we have ${\sf hw}(G)\leq f({\sf hw}^{\prime}(G))$ and ${\sf hw}^{\prime}(G)\leq f({\sf hw}(G)).$ In what follows, we denote this equivalence relation between graph parameters using the symbol $\sim$,∼ , so our result may be stated as ${\sf hw}\sim{\sf hw}^{\prime}.$ This indicates that $K_{k}$-minor-free graphs are those that are tree-decomposable into pieces which are “treewidth close” to graphs of small Euler-genus. Moreover, it restates the GMST by substituting the presence of apices and vortices [^3] See Section 3 for a formal definition of apices and vortices. with a graph modification measure: the removal of a “low-treewidth” modulator.
Dyck-grids.
Observe that the definition of ${\sf hw}^{\prime}$ in (2) is simultaneously minimizing two quantities: the treewidth of the modulator and the Euler-genus of the remaining graph. The min-max duality of the GMST is linking this entity with the exclusion of a parametric graph as a minor. In this case, the parametric graph is the complete graph on $k$ vertices.
But what about fixing the Euler-genus? What should replace cliques as the “excluded entity” to provide a min-max duality for some particular surface? In the “bottom” case of such a theorem we should position the case where the target condition is void. In this bottom case the resulting min-max theorem is the Grid Theorem asserting the equivalence between treewidth and the exclusion of grids (expressing planarity).
@@
To present our main result and give an answer to the question above, we define three parametric graphs $\mathscr{A}=\langle\mathscr{A}_{k}\rangle_{k\in\mathbb{N}},$ $\mathscr{H}=\langle\mathscr{H}_{k}\rangle_{k\in\mathbb{N}},$ and $\mathscr{C}=\langle\mathscr{C}_{k}\rangle_{k\in\mathbb{N}}$ as follows:
The annulus grid $\mathscr{A}_{k}$ is the $(4k,k)$-cylindrical grid [^4] An $(n\times m)$-cylindical grid is a Cartesian product of a cycle on $n$ vertices and a path on $m$ vertices. depicted in the left of Figure 1. The handle grid $\mathscr{H}_{k}$ (resp. crosscap grid $\mathscr{C}_{k}$) is obtained adding in $\mathscr{A}_{k}$ edges as indicated in the middle (resp. right) part of Figure 1. We refer to the added edges as transactions of the handle grid $\mathscr{H}_{k}$ or the crosscap grid $\mathscr{C}_{k}.$
Let now $h\in\mathbb{N}$ and $c\in[0,2].$ We define the parametric graph $\mathscr{D}^{h,c}=\langle\mathscr{D}_{k}^{h,c}\rangle_{k\in\mathbb{N}}$ by taking one copy of $\mathscr{A}_{k},$ $h$ copies of $\mathscr{H}_{k},$ and $c\in[0,2]$ copies of $\mathscr{C}_{k},$ then “cut” them along the dotted red line, as in Figure 1, and join them together in the cyclic order $\mathscr{A}_{k},\mathscr{H}_{k},\ldots,\mathscr{H}_{k},\mathscr{C}_{k},\ldots,\mathscr{C}_{k},$ as indicated in Figure 2. We call the graph $\mathscr{D}_{k}^{(h,c)}$ the Dyck-grid of order $k$ with $h$ handles and $c$ crosscaps. For technical reasons, we make the convention that $\mathscr{D}^{(-1,2)}_{k}=\mathscr{D}^{(0,0)}_{k}.$
Let $g\in\mathbb{N}\cup\{-1\}$ be an integer. If $g\geq 0,$ we define $\mathcal{E}_{g}$ as the class of graphs that are embeddable in some surface $\Sigma$ of Euler-genus at most $g.$ We also set $\mathcal{E}_{-1}$ to be the class containing only the empty graph and we use $\mathcal{G}_{\text{all}}$ to denote the class of all graphs. We define the graph parameter $g\mbox{-}{\sf tw}\colon\mathcal{G_{\text{all}}}\to\mathbb{N}$ as follows
@@
We stress that the only difference between (2) and (3) is that, in ($,$ the target property is fixed to $\mathcal{E}_{g}.$ Notice that $g\mbox{-}{\sf tw}(G)$ can be seen as a parametric extension of surface embeddability in the sense that if $g\mbox{-}{\sf tw}(G)\leq k$, then $G$ can be decomposed to graphs that are “$k$-treewidth close” to being embeddable in a surface of Euler-genus $g$. Observe that, for every graph $G,$ ${\sf tw}(G,V(G))={\sf tw}(G).$ Therefore, in the base case where $g=-1,$ we have that $(-1)\mbox{-}{\sf tw}={\sf tw}$.
Our next step is to define the graph parameter ${g}\mbox{-}\mathsf{bg}$ for $g\in\mathbb{N}$ as follows
${g}\mbox{-}\mathsf{bg}(G)\coloneqq\begin{cases}\max\{k\mid\mathscr{D}^{(0,0)}_{k}\leq G\}&\mbox{if $g=0$}\\ \max\{k\mid\mathscr{D}^{(h+1,0)}_{k}\leq G\mbox{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\mathscr{D}^{(h,1)}_{k}\leq G\}&\mbox{if $g=2h+1\geq 1$}\\ \max\{k\mid\mathscr{D}^{(h+1,0)}_{k}\leq G\mbox{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\mathscr{D}^{(h,2)}_{k}\leq G\}&\mbox{if $g=2h+2\geq 2$}\end{cases}$ (4)
Notice that, in case $g=0$, ${g}\mbox{-}\mathsf{bg}(G)$ is the maximum $k$ for which $G$ contains the annulus $\mathscr{A}_{k}$ as a minor. This is asymptotically equivalent to the size of the biggest grid minor of $G$. This motivated the choice of $\mathsf{bg}$ for the name of this parameter.
We are now ready to state our main result that is the following min-max equivalence between the parameters defined in (3) and (4).
Theorem 1.1.
For every $g\in\mathbb{N}\cup\{-1\},$ the parameters ${g}\mbox{-}\mathsf{tw}$ and ${(g+1)}\mbox{-}\mathsf{bg}$ are asymptotically equivalent. That is, ${g}\mbox{-}\mathsf{tw}\sim{(g+1)}\mbox{-}\mathsf{bg}.$
One important contribution of Theorem 1.1 to the general theory of graph minors is, that it makes the dependencies on the size $|V(H)|$ of the excluded minor $H$, and its Euler-genus visible and utilizable. As a consequence of our combinatorial results we are able to decuce a number of improvements on deep algorithmic results in the realm of $H$-minor-free graphs. In particular, Theorem 1.1 allows to replace the dependency on the size of $H$ in the multiplicative constants of the linear grid theorem in $H$-minor-free graphs [29], the constant-factor polynomial-time approximation algorithm for treewidth of $H$-minor-free graphs [8], and the balanced separator theorem for $H$-minor-free graphs [30], by the Euler-genus of $H$.
1.2 Surfaces and their obstructions
Theorem 1.1 is formulated in terms of the Euler-genus, so it does not give an answer to our initial question in (1) on the parameter $\mathsf{p}_{\Sigma}.$ However, it follows from a recent result of Gavoille and Hilaire [23] (Subsection 1.2) that Dyck-grids are exactly the minor-universal graphs for graphs embeddable in their respective surfaces. Hence, if a graph contains a large enough Dyck-grid for some surface $\Sigma$ as a minor, it also contains every graph embeddable in $\Sigma$ up to a fixed number of vertices as a minor. In fact, we obtain a much more general result that allows us to answer this question precisely. To explain our result in full generality, we shift our focus to (families of) surfaces.
Surfaces.
Given a pair $(h,c)\in\mathbb{N}\times[0,2]$ we define $\Sigma^{(h,c)}$ to be the two-dimensional surface without boundary created from the sphere by adding $h$ handles and $c$ crosscaps. If $c=0,$ the surface $\Sigma^{(h,c)}$ is an orientable surface, otherwise it is a non-orientable surface. By Dyck’s theorem [16, 21], two crosscaps are equivalent to a handle in the presence of a (third) crosscap. This implies that the notation $\Sigma^{(h,c)}$ is sufficient to denote all two-dimensional surfaces without boundary. The Euler-genus, denoted by ${\sf eg}(\Sigma),$ of the surface $\Sigma=\Sigma^{(h,c)}$ is $2h+c.$ Notice that if a surface has odd Euler-genus $g=2t+1$ then it is always homeomorphic to the non-orientable surface $\Sigma^{(t,1)},$ while there are two surfaces, up to homeomorphism, of even, say $2t,$ Euler-genus: the orientable one, namely $\Sigma^{(t,0)},$ and the non-orientable one, that is $\Sigma^{(t-1,2)}.$
We say that a surface $\Sigma$ is contained in some surface $\Sigma^{\prime}$ if $\Sigma^{\prime}$ can be obtained from $\Sigma$ by adding handles and crosscaps. We denote this containment relation by $\Sigma^{\prime}\preceq\Sigma.$ We use $\Sigma^{\varnothing}$ for the empty surface and, as a convention, we agree that $\Sigma^{\varnothing}$ is contained in all other surfaces. Clearly, only the empty graph is embeddable in $\Sigma^{\varnothing}.$ We also fix the Euler-genus of the empty surface $\Sigma^{\varnothing}$ to be $-1.$
Let $\mathbb{S}$ be a set of surfaces. We say that $\mathbb{S}$ is closed under containment, or simply closed, if for every $\Sigma\in\mathbb{S},$ every surface contained in $\Sigma$ also belongs to $\mathbb{S}.$ The correspondence between surfaces and Dyck graphs is given by the following theorem of Gavoille and Hilaire.

### v4

@@ Abstract
The Graph Minors Structure Theorem (GMST) of Robertson and Seymour states that for every graph $H,$ any $H$-minor-free graph $G$ has a tree-decomposition of bounded adhesion such that the torso of every bag embeds in a surface $\Sigma$ where $H$ does not embed after removing a small number of apex vertices and confining some vertices into a bounded number of bounded depth vortices. However, the functions involved in the original form of this statement were not explicit. In an enormous effort Kawarabayashi, Thomas, and Wollan proved a similar statement with explicit (and single-exponential in $|V(H)|$) bounds. However, their proof replaces the statement “a surface where $H$ does not embed” with “a surface of Euler-genus in $\mathcal{O}(|H|^{2})$”. In this paper we close this gap and prove that the bounds of Kawarabayashi, Thomas, and Wollan can be achieved with a tight bound on the Euler-genus. Moreover, we provide a more refined version of the GMST focussed exclusively on excluding, instead of a single graph, grid-like graphs that are minor-universal for a given set of surfaces. This allows us to give a description, in the style of Robertson and Seymour, of graphs excluding a graph of fixed Euler-genus as a minor, rather than focussing on the size of the graph.
Keywords: Graph minor, Surface embedding, Euler Genus, Graph structure theorem.
@@ 1 Introduction
A graph $H$ is a minor of a graph $G$ if it can be obtained from $G$ by a sequence of vertex deletions, edge deletions, and edge contractions. In their seminal series of papers on graph minors, Robertson and Seymour (R&S) [23] proved a general structure theorem for graphs that exclude a fixed graph $H$ as a minor. To state this theorem, we need a couple of definitions. Let $a$ and $b$ be non-negative integers. We say that a graph $G$ has an $a$-almost embedding in a surface $\Sigma$ of breadth $b$ if there exists a set $A\subseteq V(G),$ called the apex set, such that $G=G_{0}\cup G_{1}\cup\dots\cup G_{b}$ where
@@
    $G_{0}$ can be drawn on $\Sigma$ without crossings such that there are $b$ distinct faces [^1] These faces are called the vortices of $G.$ See Subsection 4.3 for a formal definition and Subsection 4.3 specifically for a definition of the path decomposition (we call it a “linear decomposition” later on) for vortices. $F_{1},\dots,F_{b}$ and for every $i\in[b]$ the vertices of $V(G_{0})\cap V(G_{i})$ all lie on $F_{i},$ and
@@
    for each $i\in\{1,\dots,b\},$ $G_{i}$ is a graph with a path decomposition [^2] For the definition of tree decompositions see Section 2. A path decomposition of a graph $G$ is a tree decomposition $(T,\beta)$ for $G$ where $T$ is a path. of width at most $b$ such that the vertices of $V(G_{0})\cap V(G_{i})$ appear in distinct bags of this decomposition exactly in the order as they appear on $F_{i}.$
A graph $G$ is a clique sum of two graphs $G_{1}$ and $G_{2}$ if there exist cliques $S_{i}\subseteq G_{i}$ for both $i$ such that $|S_{1}|=|S_{2}|$ and $G$ can be obtained from $G_{1}$ and $G_{2}$ by identifying the vertices of $S_{1}$ with those of $S_{2}$ into a clique $S$ and possibly removing some of the edges with both endpoints in $S.$
With these definition, the Graph Minor Structure Theorem of Robertson and Seymour reads as follows.
[…]
There exists a function $f\colon\mathbb{N}\to\mathbb{N}$ such that for all graphs $H$ and $G$ either
  1. 1.
    $H$ is a minor of $G,$ or
  2. 2.
    $G$ can be obtained by means of clique sums from graphs that are $f(|V(H)|)$-almost embeddable with breadth at most $|V(H)|^{2}$ in surfaces where $H$ does not embed.
Robertson and Seymour did not provide any bounds on the function $f.$ An important lower bound on $f(|V(H)|)$ is the treewidth an $H$-minor-free graph should have such that the methods of R&S can be applied. Moreover, a particularly difficult to handle contributions to its upper bound stems from the fact that, in order to prove the “where $H$ does not embed”-part of Section 1, one needs to show that there is some function $g$ such that for every $\mathsf{h},$ any graph that embeds on a surface $\Sigma$ with representativity [^3] Roughly: the minimum length of a non-contractible cycle. at least $g(h)$ will contain all graphs on $\mathsf{h}$ vertices that embed in $\Sigma$ as a minor (see Theorem 4.3 in [23]). This proof, however, is non-constructive.
In recent years, many steps at making the graph minors theory of R&S constructive have been made. Arguably the biggest step here is a new, and fully constructive, proof of a variant of Section 1 by Kawarabayashi, Thomas, and Wollan. Their version of the GMST reads as follows.
[…]
There exists a function $f\colon\mathbb{N}\to\mathbb{N}$ with $f(k)\in 2^{\mathsf{poly}(k)}$ such that for all graphs $H$ and $G$ either
  1. 1.
    $H$ is a minor of $G,$ or
  2. 2.
    $G$ can be obtained by means of clique sums from graphs that are $f(|V(H)|)$-almost embeddable with breadth at most $|V(H)|^{2}$ in surfaces of Euler genus $\mathcal{O}(|V(H)|^{2})$.
Notice that the bound on the Euler genus of the surfaces involved in Section 1 is significantly worse than the bound given in Section 1 where it is optimal. The reason for this worse bound is the non-constructiveness in the original proof by R&S [23] and the determination of Kawarabayashi, Thomas, and Wollan to avoid these arguments in order to make their theorem constructive.
Local structure theorems.
Both, Section 1 and Section 1 are theorems that describe the structure of $H$-minor-free graphs globally. However, both theorems, as most results in this area, are based on much more technical local versions. The world “local” refers to the structure in an $H$-minor-free graph that is governed by a large wall [^4] See Subsection 4.1 for related definitions of walls and tangles or, in other words, the structure that is controlled by a large-order tangle. Let $W$ be an $r$-wall for some $r\geq 3.$ A simplification of $W$ is an $r$-wall $W^{\prime}$ which is obtained by contracting some edges of $W.$ In this language one can restate Section 1 in its local form as follows.
[…]
There exists functions $f,g\colon\mathbb{N}\to\mathbb{N}$ with $f(k),g(k)\in 2^{\mathsf{poly}(k)}$ such that for all graphs $H$ and all graphs $G$ containing a $g(|V(H)|)$-wall $W$ either
  1. 1.
    there is a minor-model of $H$ which is highly connected [^5] The term “highly connected” here is slightly informal. Formally this means that there exists a minor model $\{X_{v}\}_{v\in V(H)}$ of $H$ such that each $X_{v}$ intersects many intersections of horizontal and vertical paths of $W.$ to $W,$ or
  2. 2.
    $G$ can be obtained by means of clique sums from graphs that are $H$-minor-free and a graph $G_{W}$ which is $f(|V(H)|)$-almost embeddable with breadth at most $|V(H)|^{2}$ in a surface of Euler genus $\mathcal{O}(|V(H)|^{2})$ such that there is a simplification $W^{\prime}$ of $W$ with $W^{\prime}\subseteq G_{W}.$
The local structure theorem describes a single summand in the clique-sum structure provided by statements like Section 1 and Section 1. While such a local structure theorem only describes those summands of large treewidth, this is usually sufficient as all summands of small treewidth can be further split into summands of small size (which is the essence of the definition of treewidth). A simple greedy procedure, first introduced by R&S [22], then allows to immediately go from a local structure theorem to a global one [^6] In case of this paper, such a proof can be found in Section 7. while increasing the involved function at most by some polynomial.
@@ 1.1 Our contribution
Recently, Gavoille and Hilaire [13] proved the existence of a minor-universal graph for all graphs that embed in a fixed surface $\Sigma.$ In particular, the bounds they achieve are polynomial in the Euler genus of $\Sigma,$ that is, they prove that for every integer $\mathsf{h},$ their universal graph of order $\mathsf{poly}(g\cdot h)$ contains every graph on at most $\mathsf{h}$ vertices that embeds in $\Sigma$ as a minor. Here $g$ denotes the Euler genus of $\Sigma.$ We leverage this result to prove the following local version of the GMST.
Theorem 1.1.
There exists functions $f,g\colon\mathbb{N}\to\mathbb{N}$ with $f(k),g(k)\in 2^{\mathsf{poly}(k)}$ such that for all graphs $H$ and all graphs $G$ containing a $g(|V(H)|)$-wall $W,$ either
  1. 1.
    $G$ contains a minor-model of $H$ which is highly connected to $W,$ or
  2. 2.
    $G$ can be obtained by means of clique sums from graphs that are $H$-minor-free and a graph $G_{W}$ which is $f(|V(H)|)$-almost embeddable with breadth at most $|V(H)|^{2}$ in a surface where $H$ does not embed such that there is a simplification $W^{\prime}$ of $W$ with $W^{\prime}\subseteq G_{W}.$
With this, we achieve a first unification of the tight bound on the Euler genus from Section 1 and the constructive bound on the function $f$ in Section 1.
This is done in two steps. First, we define several families of grid-like graphs representing a fixed surface $\Sigma,$ then we show in Section 3 that, up to a function depending linearly on the order of the graph and exponentially on the Euler genus of $\Sigma,$ all of these graphs are equivalent. See Figure 1 for some examples of such graphs.
@@
This allows us to apply a recent result of Gavoille and Hilaire [13] to show that there is some universal constant $c_{\Sigma}$ such that any of these grid-like graphs of order $c_{\Sigma}t^{2}$ contains every graph on at most $t$ vertices that embeds in $\Sigma$ as a minor. This means, we leverage the theorem of Gavoille and Hilaire to prove that their minor-universal graph exists for any given “signature” (see [13] for a definition). This is important for the following reason. We know that any surface $\Sigma$ can be obtained from the sphere by adding $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps where the number $2\mathsf{h}+\mathsf{c}$ is precisely the Euler genus of $\Sigma.$ By the Classification Theorem of Surfaces, also known as Dyck’s Theorem, [9, 12], we know that every surface is homeomorphic to a surface that is obtained from the sphere by adding $\mathsf{h}$ handles and at at most two crosscaps. Gavoille and Hilaire proved that there exists a minor-universal graph for orientable surfaces that is obtained by only adding handles to the sphere, and one for non-orientable surfaces obtained by only adding crosscaps to the sphere. However, the proof of Section 1 by Kawarabayashi, Thomas, and Wollan, while guaranteeing a bound on their total number, extracts some number of handles and crosscaps in arbitrary order.
Let us give a quick definition of these grid-like graphs. The annulus grid $\mathscr{A}_{k}$ is the $(4k,k)$-cylindrical grid [^7] An $(n\times m)$-cylindrical grid is a Cartesian product of a cycle on $n$ vertices and a path on $m$ vertices. depicted in the left of Figure 1. The handle grid $\mathscr{H}_{k}$ (resp. crosscap grid $\mathscr{C}_{k}$) is obtained adding in $\mathscr{A}_{k}$ edges as indicated in the middle (resp. right) part of Figure 1. We refer to the added edges as transactions of the handle grid $\mathscr{H}_{k}$ or the crosscap grid $\mathscr{C}_{k}.$
Let now $\mathsf{h}\in\mathbb{N}$ and $\mathsf{c}\in[0,2].$ We define the graph $\mathscr{D}_{k}^{(\mathsf{h},\mathsf{c})}$ by taking one copy of $\mathscr{A}_{k},$ $\mathsf{h}$ copies of $\mathscr{H}_{k},$ and $\mathsf{c}\in[0,2]$ copies of $\mathscr{C}_{k},$ then “cut” them along the dotted red line, as in Figure 1, and join them together in the cyclic order $\mathscr{A}_{k},\mathscr{H}_{k},\ldots,\mathscr{H}_{k},\mathscr{C}_{k},\ldots,\mathscr{C}_{k},$ as indicated in Figure 2. We call the graph $\mathscr{D}_{k}^{(\mathsf{h},\mathsf{c})}$ the Dyck-grid of order $k$ with $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps. For technical reasons, we make the convention that $\mathscr{D}^{(-1,2)}_{k}=\mathscr{D}^{(0,0)}_{k}.$
Notice that, in a similar way, we may take any pair of non-negative integers $\mathsf{h}$ and $\mathsf{c}$ as input and concatenate one copy of $\mathscr{A}_{k}$ together with $\mathsf{h}$ copies of $\mathscr{H}_{k}$ and $\mathsf{c}$ copies of $\mathscr{C}_{k}$ in any order. Any such graph yields a (parametric) representation of the surface obtained from the sphere by adding $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps.
@@
We prove that any such object with $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps can be translated into any other such object with $\mathsf{h}^{\prime}$ handles and $\mathsf{c}^{\prime}$ crosscaps while preserving $2\mathsf{h}+\mathsf{c}=2\mathsf{h}^{\prime}+\mathsf{c}^{\prime}$ where $\mathsf{c}^{\prime}\neq 0$ if and only if $\mathsf{c}\neq 0$ and $\mathsf{c}^{\prime}$ is even if and only of $\mathsf{c}$ is even. One may see this result as the graph-theoretical analogue of Dyck’s Theorem.
Second, we use an intermediate structure from [17] called a $\Sigma$-configuration [^8] See Section 5 for the definition. to prove that Section 1 may be extended to provide, as a third outcome, one of the Dyck-grids of large order representing some predetermined orientable surface $\Sigma_{1}$ or some predetermined non-orientable surface $\Sigma_{2}.$ This further allows us to adjust the second outcome in a way that only allows surfaces whose Euler genus is strictly smaller than the Euler genus of $\Sigma_{1}$ and $\Sigma_{2}.$
Using this result, we do not only prove Theorem 1.1, but we also prove a variant of Section 1 which, instead of a clique, excludes exclusively two Dyck-grids, one for an orientable surface and one for a non-orientable one. This theorem reads as follows.